# Lessons on Datasets and Paradigms in Machine Learning for Symbolic Computation: A Case Study on CAD

### v1

@@ 4.1.3. Illustrative example
For illustration purposes, we will consider a hypothetical example which producues the timings to build a three-variable CAD indicated in Table 1. Figure 1 then illustrates how we create ML classification instances from our working example. This will serve as a comparison for the alternative methodologies presented later.
@@ 4.2. Imbalance in the dataset
As observed recently in [15] and [22] an analysis on the classification dataset described in Section 4.1 reveals that the dataset obtained is imbalanced: there are far more pairs of labels with ordering $\succ_{123}$ than any of the others, as can be observed in Figure 5. Such an imbalance is a problem for ML, because models trained on an imbalance dataset will be biased towards the majority class. Moreover, this problem may go unnoticed if the testing dataset is also imbalanced in a similar way, reporting misleading results that will not generalise when testing on a balance dataset.
@@ 4.3. Improving the datasets with data augmentation
Our instances consist of sets of polynomials, for example, $\{x_{1}^{2}-x_{2},x_{3}^{3}-1\}$ for which we can determine, by computing and comparing CADs, that the optimal variable ordering to compute an CAD for this set is $x_{2}\succ x_{1}\succ x_{3}$. Now observe that simply by swapping the names of the variables $x_{1}$ and $x_{2}$ we may obtain the new set of polynomials $\{x_{2}^{2}-x_{1},x_{3}^{3}-1\}$, in which we know, without any further CAD computation, that the optimal variable ordering is $x_{1}\succ x_{2}\succ x_{3}$.
@@
Figure 2. Data augmentation by swapping $x_{1}$ and $x_{2}$ for the illustrative example
@@ 4.3.1. Three datasets
We will refer to the dataset described in Section 3.1 as the “imbalanced dataset”. Similar to how the arrow image dataset can be balanced by randomly rotating, our dataset can be balanced by randomly permuting the variable names: we refer to the dataset obtained this was as the the “balanced dataset”. We then go further and fully augment the dataset by including all possible variable permutations, obtaining six instances from each original one: creating the “augmented dataset”. The distribution of instances to optimal orderings is illustrated in Figure 5 -- Figure 5.
@@ 4.3.2. Other ideas to augment the dataset
The equivalent of zooming an image for a set of polynomials could be making changes of variables of the form $y_{i}=j_{i}x_{i}+k_{i}$, where $j_{i},k_{i}\in\mathbb{R}$ for all $i$. These changes of variables do affect the features used to describe the set of polynomials, mainly because they tend to change the sparsity of the features, one of the values reflected by the features used, so they would definitely create new instances. However, these changes in sparsity also affect the computations of resultants and discriminants needed to build a CAD, so the label of the instance may change; meaning that constructing CADs for all possible orderings would be necessary to find the real label of the instance.
@@ 4.4. Replicating prior work
Figure 6. Existing heuristics and ML models trained on the biased dataset, evaluated over the biased dataset.
@@ 4.5. ML classifiers obtained when balancing and augmenting the training dataset
So models trained imbalanced dataset perform poorly on a balanced one. We will test whether using the balanced dataset for training the models will improve the performance: this should be the case since the models will no longer learn the biases they got when trained in an imbalanced dataset. We will also evaluate if using the augmented dataset described in Section 4.3 offer further benefits: it has six times the number of instances of the imbalanced and balanced datasets.
@@
Training dataset Total time Completed Accuracy Markup
@@
Training dataset Total time Completed Accuracy Markup
@@
Training dataset Total time Completed Accuracy Markup
@@
Training dataset Total time Completed Accuracy Markup

### v2

@@ 4.1.3. Illustrative example
For illustration purposes, we will consider a hypothetical example that produces the timings to build a three-variable CAD indicated in Table 1. Figure 2 then illustrates how we create ML classification instances from our working example. This will serve as a comparison for the alternative methodologies presented later.
@@ 4.2. Imbalance in the dataset
As observed recently in [16] and [23] an analysis on the classification dataset described in Section 4.1 reveals that the dataset obtained is imbalanced: there are far more pairs of labels with ordering $\succ_{123}$ than any of the others, as can be observed in Figure 6. Such an imbalance is a problem for ML, because models trained on an imbalanced dataset will be biased towards the majority class. Moreover, this problem may go unnoticed if the testing dataset is also imbalanced in a similar way, reporting misleading results that will not generalise when testing on a balanced dataset.
@@ 4.3. Improving the datasets with data augmentation
Our instances consist of sets of polynomials, for example, $\{x_{1}^{2}-x_{2},x_{3}^{3}-1\}$ for which we can determine, by computing and comparing CADs, that the optimal variable ordering to compute a CAD for this set is $x_{2}\succ x_{1}\succ x_{3}$. Now observe that simply by swapping the names of the variables $x_{1}$ and $x_{2}$ we may obtain the new set of polynomials $\{x_{2}^{2}-x_{1},x_{3}^{3}-1\}$, in which we know, without any further CAD computation, that the optimal variable ordering is $x_{1}\succ x_{2}\succ x_{3}$.
@@
Figure 3. Classification training workflow with data augmentation: $S^{\prime}$ has been created by renaming $x_{1}$ to $y_{2}$, $x_{2}$ to $y_{1}$ and $x_{3}$ to $y_{3}$ in $S$.
@@ 4.3.1. Three datasets
We will refer to the dataset described in Section 3.1 as the “imbalanced dataset”. Similar to how the arrow image dataset can be balanced by randomly rotating, our dataset can be balanced by randomly permuting the variable names: we refer to the dataset obtained this was as the “balanced dataset”. We then go further and fully augment the dataset by including all possible variable permutations, obtaining six instances from each original one: creating the “augmented dataset”. The distribution of instances to optimal orderings is illustrated in Figure 6 -- Figure 6.
@@ 4.3.2. Other ideas to augment the dataset
The equivalent of zooming an image for a set of polynomials could be making changes of variables of the form $y_{i}=j_{i}x_{i}+k_{i}$, where $j_{i},k_{i}\in\mathbb{R}$ for all $i$. These changes of variables do affect the features used to describe the set of polynomials, mainly because they tend to change the sparsity of the variables, one of the features used, so they would definitely create new instances. However, these changes in sparsity also affect the computations of resultants and discriminants needed to build a CAD, so the label of the instance may change; meaning that constructing CADs for all possible orderings would be necessary to find the real label of the instance.
@@ 4.4. Replicating prior work
Figure 7. Existing heuristics and ML models trained on the biased dataset, evaluated over the biased dataset. See Section 2.3 for the definitions of gmods, brown and T1, and see Section 3.6 for definitions of the other acronyms.
@@ 4.5. ML classifiers obtained when balancing and augmenting the training dataset
So models trained imbalanced dataset perform poorly on a balanced one. We will test whether using the balanced dataset for training the models will improve the performance: this should be the case since the models will no longer learn the biases they got when trained in an imbalanced dataset. We will also evaluate if using the augmented dataset described in Section 4.3 offers further benefits: it has six times the number of instances of the imbalanced and balanced datasets.
@@
Training dataset Total time Completed Accuracy Markup
[…]
@@
Training dataset Total time Completed Accuracy Markup
[…]
@@
Training dataset Total time Completed Accuracy Markup
[…]
@@
Training dataset Total time Completed Accuracy Markup
[…]
@@
5. Changing from the Classification to the Regression Paradigm
As we described in Section 1.1, although prior work on ML to make choices for our problem has employed the classification paradigm, since the choices seek to minimise a continuous metric the problem may also be conceived as ML regression. We explore the effects of this paradigm change in this section: first describing how regression instances are created in Section 5.1, then presenting an experimental comparison in Section 5.2.
5.1. Creating a regression instance
A regression instance consists of a pair: an input and a target. In our case, the input is formed from the set of polynomials and variable ordering, and the target will be the time taken to build a CAD for these. The embedding that we will use to encode the input will be formed by the same three vectors of features introduced previously in Section 4.1.1 for classification. However here, as illustrated in Figure 9, the order in which we append the vectors will be determined by the input variable ordering.
Figure 9. Regression training workflow showing how a single regression instance is created from just 1 of the 6 orderings for the example of Section 4.1.3; 5 further instances may be created similarly.
This embedding eliminates all existing symmetries in the classification embedding, so it is not possible to augment this dataset as we did in Section 4.3 for classification. However, the number of instances in the regression dataset is already as large as that of the augmented classification dataset since each classification instance becomes six regression instances. Furthermore, the regression dataset encodes more information than the augmented classification dataset. Although instances in both datasets share the same features, their labels differ: the augmented classification dataset labels are binary, indicating only which ordering is the fastest, while the regression dataset labels contain numerical information about the exact time each ordering took.
Our ML regression models will predict how long it takes to build a CAD for a given set of polynomials using a given variable ordering. However, this is not our final goal: we are interested in choosing variable orderings to build a CAD for a given set of polynomials as fast as possible. Thus, for a given set of polynomials in the testing dataset, we must ask the regression model to estimate how long it would take to build a CAD for each of the variable orderings, and then we will choose the ordering that was estimated to take the shortest time, as illustrated in Figure 10. This final choice can then be evaluated using the same metrics introduced in Section 3.7.
Figure 10. Overview of how the regression model is used.
5.2. Results comparing classification with regression
Figure 11 compares the total time required to build the CADs in our test dataset using the variable ordering suggested by each model. We see that the KNN, and SVM, models all perform better under the regression paradigm, while the RF, MLP, and XGBoost models perform worse. Overall, the best performance on the dataset is achieved by a model under the regression paradigm. So the results partially validate our hypothesis that regression can use extra information to improve performance on the choice, but with the substantial caveat that it is model dependent.
Figure 11. Comparing the total times to build CADs for the balanced dataset with the variable ordering suggested by classifiers (Class) and regressors (Reg)
We also reflect that the scope of choices that can be made using regression is broader than those that can be made using classification. For example, when using CAD for quantifier elimination, we must project variables in the order of quantification but still have freedom to change order within quantification blocks (and the free variables). In this case, if a classifier returns an option that does not satisfy these restrictions the recommendation is useless; however, using the regression paradigm, one can simply take the best choice that meets the restrictions.
5.3. Limitations with respect to number of variables
The number of possible outputs the classification models have to choose from is factorial of the number of variables. A classification model trained for $n$ variables cannot study a problem with more variables; and although the methodology could generalise (i.e., we may train new classifiers with a dataset of problems in more variables), the factorial growth means this will get harder quickly.
In comparison the regression models presented in this section always produced a single real number regardless of the number of variables. The difficulty in generalising to more variables will be the embedding used to input sets of polynomials into both the regression and classification models: this will still grow with the number of variables, but only linearly (27 times the number of variables).
If we want to choose variable orderings for an arbitrary number of variables, then ideally we should find an embedding that has an invariant size with respect to the number of variables. We consider one approach to this in Section 6, in which our model takes always a fixed-size input of only 27 features.
6. Using Regression to Rank Variables
In Section 5 we trained regression models to estimate how long it would take to build a CAD for a problem with a given variable ordering. In this section we will use ML to rank the variables themselves instead of the variable orderings, allowing for an easier generalisation to more variables.
This is similar to how the human-designed heuristics in Section 2.3 choose a variable ordering. Human heuristics assign some measure of complexity to the variables and pick the simpler variables first in the ordering, but their measure of complexity is defined by the creator of the heuristic and, therefore, is fixed. Instead, we will train ML models so that they learn their own measures of complexity, and we will pick the variables that the model considers to be simpler first in the ordering. This shares a similarity with Chen et al. (2020) [12] who also proposed to train models to choose variables instead of orderings. However, they still trained classification models and thus suffer from the limitations discussed above, being limited in their approach to sets of polynomials with less than ten variables.
Moreover, instead of choosing the whole variable ordering using only the original polynomials, once the first variable has been chosen, we will use the projected set to choose the next variable. This allows us to make our choices using the most recent information.
6.1. Creating instances for regression on variables
Our instances now concern the polynomials and a single variable. For an embedding we use just the features for that variable, i.e. $f_{i}(S)$: the features describing variable $x_{i}$ in a set of polynomials $S$.
In Section 5 we trained ML models to learn the complexity of a variable ordering and there was a clear output to use: the time necessary to build a CAD using such variable ordering. In this section, for estimating the complexity of a variable, we have available the timings of the orderings that start with that variable: in our three-variable case study there are two possible orderings that begin with each variable. We use the quickest of these to label the instance, as illustrated in Figure 12.
Figure 12. Illustration of how the instance for variable $x_{1}$ is encoded for the example of Section 4.1.3. Note that two further instances for the other variables will be created similarly.
6.2. Results
Tables 6.2 -- 6.2 compare the classification paradigm (trained with the augmented dataset), the regression paradigm on the orderings, and the regression paradigm on the variables. Interestingly, we see that the two models which did poorest under the regression paradigm to pick an ordering MLP and XGBoost do better under this alternative regression paradigm (while the others do worse). The lowest timing overall remains the KNN under the previous regression paradigm, but given the wider applicability of this alternative paradigm we think it worthy of consideration.
In Figure 13 we show the spread of timings for all the key strategies presented in this paper, represented by letters as below. They show that for all models the original approach of training on an unbalanced dataset does not generalise to a balanced dataset (A vs B); but that this performance is regained and sometimes exceeded by the new approaches of this paper (C -- F). The optimal paradigm does seem to depend on the ML model type being used.
Short name Training dataset Testing dataset Paradigm
A Imbalanced (see Section 4.2) Imbalanced Classification
B Imbalanced Balanced Classification
C Balanced (see Section 4.3) Balanced Classification
D Augmented (see Section 4.3) Balanced Classification
E Regression orderings (See Section 5.1) Balanced Regression
F Regression variables (see Section 6) Balanced Regression
Strategy Total time Completed Accuracy Markup
@@ F Regression variables (see Section 6) Balanced Regression
Table 7. Metrics for strategies using a KNN model over the balanced dataset.
Strategy Total time Completed Accuracy Markup
@@
Table 8. Metrics for strategies using a RF model over the balanced dataset.
Strategy Total time Completed Accuracy Markup
@@
Table 9. Metrics for strategies using an SVM model over the balanced dataset.
Strategy Total time Completed Accuracy Markup
@@
Table 10. Metrics for strategies using a MLP model over the balanced dataset.
Strategy Total time Completed Accuracy Markup
[…]
Table 11. Metrics for strategies using an XGB model over the balanced dataset.
(a) KNN models.
(b) RF models.
(c) SVM models.
(d) MLP models.
(e) XGBoost models.
Figure 13. The total times achieved by models on the balanced testing dataset under different strategies
7. Final Thoughts
7.1. Conclusions on data balance and augmentation
7.1.1. The importance of data exploration
We found in Section 4 that this well known dataset was imbalanced for our problem and that if we trained classifiers with such imbalanced data they could not perform well on balanced data. This demonstrates the importance of exploring the data before moving to ML. Imbalance is not always inappropriate for ML: some applications will naturally have imbalanced data, and ML models should be aware of this. However, in our case, we seek heuristics for choosing variable orderings for CAD applied in general, and there is little rationale to suppose that general CAD applications favour one ordering over another444except perhaps the existence of the SMT-LIB data!. Thus, our advice is to ensure that ML models for such applications are trained on balanced data.
7.1.2. Recovering performance through balancing and augmentation
The good news is that ML performance can be improved simply by training on balanced data, reducing the total time by 28% on average, re-validating the value of the data science-led approach. A key conclusion here is that it is beneficial to go further and use maximum data augmentation: all models benefited from this beyond just balancing the data, no matter which dataset they are tested on. Training in an augmented dataset of six times the size of the original dataset reflected in an increase in accuracy of 40% on average and a reduction of 38% in the total time on average. In fact, the performance lost from the original imbalanced case is almost fully recovered this way.
7.1.3. Comparison with the work of Hester et al. (2023)
Some of these conclusions were also drawn in a recent paper [23]. Table 2 in [23] presents the accuracies of trained models on different datasets: note that both their “Training Set 2” and “Dataset 1” contain instances in which the models have been trained, meaning that the former is the most appropriate column for evaluation in that table: that column shows results similar to ours.
We note that the original dataset in [23] contained 6895 instances, while in our paper the initial dataset only contained 1019 instances. This is because, even though both datasets have the same ultimate source (the SMT-LIB), our dataset had been stripped of potential duplicate instances (sets of polynomials whose CAD tree structures are potentially identical for every variable ordering), as described in detail in Section 3.1. We view this as a necessary step to make meaningful use of the QF_NRA section of the SMT-LIB where there are many very similar problems.
This comparison with [23] shows that the size alone of the dataset is not what matters (since [23] has a similar accuracy to the models presented here despite training with much more data). Rather, it is the number of qualitatively different problems within the dataset. That is, there is little benefit to including multiple very similar problems. It may seem that data augmentation adds no new information, but since the ML models are not aware of these symmetries by exposing them with augmentation, we actually give them access to this information.
7.2. Conclusions on ML paradigms
Despite the fact that the regression dataset contains more information than the classification one, such as the specific timings of each ordering, this did not always lead to improvement in the performance of regression models compared to classification models: it seemed that some models were better able to do this than others and this choice of paradigm should be optimised along with model and hyperparameter selection in the ML workflow.
The overall lowest timings were achieved by one of the new regression approaches. The differences in performance were not huge: as argued in Section 5.2, we suggest that more important than performance statistics is the wider scope of choices that can be made in symbolic computation using the new regression methodologies. Consider for example an application to $S$-pair selection in Buchberger’s algorithm: we do not know in advance of a decision how many pairs there will be to choose from, but this may be tackled using the paradigm presented in Section 6, and still under a supervised learning paradigm rather than using reinforcement learning as in [33].
7.3. Future work
A key area for future work could be the embeddings: those we use still ignore much information from the polynomials such as all coefficient information. Jia et al. (2023) [27] already proposed a graph embedding that allows inputting a lot of information about a set of polynomials into a GNN model, and they experience a performance similar to that of the heuristic gmods in the SMT-LIB dataset and even beat this heuristic in a randomly generated dataset.
Given the success of this data augmentation, an obvious area for future work is to look for additional augmentation techniques. Returning to the computer vision analogy: rotations are not the only augmentation tool; there are others also, e.g. mirror reflections. Regarding mathematical objects, a corresponding augmentation technique may be substituting a variable with its negative, which would create a new instance without the need for any further labelling. We could also consider more involved variable transformations; however, these would most likely require additional CAD computations for data labelling, which is the most expensive part of this whole process.
Finally, we note that all these ideas would generalise easily to variable ordering choice for the other decision procedures of non-linear real arithmetic commonly found in the wider toolchains of the SC2 community. Further, the lessons we illustrated on the importance of exploratory data analysis and paradigm choice would be useful to many other choices available in symbolic computation.
Data Access Statement
This work in this paper formed part of Tereso del Río’s PhD thesis. All code used for the experiments reported on in this paper (including to generate the results images ) can be found within the data release for that thesis here: https://zenodo.org/doi/10.5281/zenodo.10834972.
Acknowledgments
Tereso del Río is supported by Coventry University and a travel grant from the London Mathematical Society (LMS). Matthew England is supported by UKRI EPSRC Grant EP/T015748/1, Pushing Back the Doubly-Exponential Wall of Cylindrical Algebraic Decomposition (the DEWCAD Project).